Acknowledgements

The authors would like to thank G. Gallavotti for helpfull discussions and one of the authors (P.W.) would like to thank T. Spencer for his continued support and A. Sokal for his enthusiasm and for all the usefull comments.

References

[ 1] J.–P. Eckmann, H. Koch, P. Wittwer. A computer–assisted proof of universality for area–preserving maps. Memoirs AMS \underline{47}, 289 (1984). [ 2] J.–P. Eckmann, P. Wittwer. Computer methods and Borel summability applied to Feigenbaum’s equation. Lecture Notes in Physics, Springer–Verlag, Berlin, Heidelberg, New York, Tokyo (1985). [ 3] O.E. Lanford III . A computer–assisted proof of the Feigenbaum conjectures. Bull. AMS, New Series \underline{6}, 127 (1984). [ 4] R.E. Moore. Interval analysis. Prentice–Hall, series in automatic computation (1966). [ 5] R.E. Moore. Methods and applications of interval analysis. SIAM, Philadelphia (1979). [ 6] A. Sokal. An improvement of Watson’s theorem on Borel summability. J. Math. Phys \underline{21}, 261 (1980). [ 7] R. de la Llave, O.E. Lanford III . Work in progress. [ 8] J.–P. Eckmann, J. Magnen, R. Sénéor. Decay properties and Borel summability for the Schwinger functions in P (ϕ)_{2} theories. Comm. Math. Phys. \underline{39}, 251 (1975). [ 9] J. Magnen, R. Sénéor. Phase space cell expansion and Borel summability for the Euclidean ϕ_{3}^{4} theory. Comm. Math. Phys. \underline{56}, 237 (1977).

[10] P. Collet, J.–P. Eckmann. A renormalization group analysis of the hierarchical model in statistical physics. Lecture Notes in Physics \underline{74}, Springer–Verlag, Berlin, Heidelberg, New York (1978).

[11] K. Gawedzki, A. Kupiainen. Triviality of ϕ_{4}^{4} and all that in a hierarchical model approximation. Journal of Statistical Physics \underline{29}, 683 (1982).

[12] K. Gawedzki, A. Kupiainen. Rigorous renormalization group and asymptotic freedom. In: Scaling and self–similarity in physics (renormalization in statistical mechanics and dynamics). Jürg Fröhlich (Ed.). Progress in Physics \underline{7}, Birkhäuser, Boston, Basel, Stuttgart (1983).

[13] P.M. Bleher. Usp. Mat. Nauk. \underline{32}, 243 (1977).

[14] F.J. Dyson. Existence of a phase transition in a one–dimensional Ising ferromagnet. Comm. Math. Phys. \underline{12}, 91-107 (1969).

[15] F.J. Dyson. Nonexistence of spontaneous magnetization in a one–dimensional Ising ferromagnet. Comm. Math. Phys. \underline{12}, 212–215 (1969).

[16] P.M. Bleher, Ja.G. Sinai. Investigation of the critical point in models of the type of Dyson’s hierarchical model. Comm. Math. Phys. \underline{33}, 23 (1973).

[17] P.M. Bleher, Ja.G. Sinai. Critical indices for Dyson’s asymptotically hierarchical models. Comm. Math. Phys. \underline{45}, 347 (1975).

[18] K.G. Wilson, J.B. Kogut. The renormalization group and the ε expansion. Phys. Report. \underline{12C}, 75 (1974).

[19] K. Wilson. Renormalization group and critical phenomena, II phase space cell analysis of critical behavior. Phys. Rev. \underline{B4}, 3184 (1971).

[20] P. Collet, J.–P. Eckmann, B. Hirsbrunner. A numerical test of Borel summability in the ε–expansion of the hierarchical model. Physics letters \underline{71B}, 385 (1977).

[21] M.C. Reed. Functional analysis and probability theory. In: G. Velo, A. Wightman (Eds.), constructive quantum field theory. Lecture Notes in Physics \underline{25}, Springer–Verlag, Berlin, Heidelberg, New York (1976).

[22] F. Treves. Topological vector spaces, distributions and kernels. Academic Press, New York, London (1967).

[23] B.W. Kernigham, D.M. Ritchie. The C programming language. Prentice–Hall, software series (1978).

[24] D. Stevenson, IEEE Computer Society. A proposed standard for binary floating–point arithmetic. Draft. 8.0 of IEEE Task P754 (1981).

[25] G. Benfatto, M. Cassandro, G. Gallavotti, F. Nicolò, E. Olivieri, E. Presutti, E. Scacciatelli. Some probabilistic techniques in field theory. Comm. Math. Phys. \underline{71}, 95-130 (1980).

[26] G. Benfatto, M. Cassandro, G. Gallavotti, F. Nicolò, E. Olivieri, E. Presutti, E. Scacciatelli. On the ultraviolet stability in the Euclidean scalar field theories. Comm. Math. Phys. \underline{71}, 343–400 (1980).

[27] G. Baker. Analysis of hyperscaling in the Ising model by the high–temperature series method. Phys.Rev. \underline{B15}, 1553-1559 (1977).

[28] G. Baker, J.M. Kincaid. The continuous spin Ising model, g_{0} : ϕ_{4}^{4}: field theory, and the renormalization group. J. Stat. Phys. \underline{24}, 469–528 (1981).

[29] J. Rosen. The Ising model limit of ϕ^{4} lattice fields. Proc. AMS \underline{66}, 114–118 (1977).

[30] G. Caginalp. The ϕ^{4} lattice filed theory as an asymptotic expansion about the Ising limit. Ann. Phys. \underline{124}, 189-207 (1980).

[31] G. Caginalp. Thermodynamic properties of the ϕ^{4} lattice field theory near the Ising limit. Ann. Phys. \underline{126}, 500-511 (1980).

[32] F. Constantinescu. Expansion of the double–well model near the Ising model. Institut für angewandte Mathematik Johann Wolfgang Goethe Universität. Preprint (1980).

[33] F. Constantinescu, B. Stroter. The Ising limit of the double–well model. J. Math. Phys. \underline{21}, 881-890 (1980).

[34] Le Guillou, Zinn–Justin. Critical exponents for the n-vector model in three dimensions from field theory. Phys. Rev. Lett. \underline{39}, 95-98 (1977).

[35] B. Nickel. Hyperscaling and universality in three dimensions. Cargese lectures (1980).

[36] C. Bender, F. Cooper, G. Guralnik, R. Roskies, D. Sharp. Numerical computation of the renormalized effective potential in the strong coupling limit. Phys. Rev. \underline{D23}, 2976 (1981).

[37] See e.g. in: Handbook of Chemistry and Physics

[38] W.J. Camp, D.M. Saul, J.P. Van Dyke, M. Wortis. Series analysis of corrections to scaling for the spin–pair correlations of the spin–s Ising model: confluent singularities, universality and hyperscaling. Phys. Rev. \underline{B14}, 3990–4001 (1976).

[39] D. Bessis, P. Moussa, G. Turchetti. Subdominant critical indices for the ferromagnetic susceptibility of the spin–\frac{1}{2} Ising model. J. Phys. \underline{A13}, 2763–2773 (1980).

[40] P. Moussa. Subdominant critical singularities in the bcc Ising model. J. Stat. Phys. \underline{27}, 711–719 (1982).

[41] J.–H. Chen, M.E. Fisher. Unbiased estimation of corrections to scaling by partial differential approximants. Phys. Rev. Lett. \underline{48}, 630–634 (1982).

[42] J. Adler, M. Moshe, V. Privman. Unbiased map of the temperature plane and its consequences. Phys. Rev. \underline{B26}, 3958–3959 (1982).

[43] M. Ferer, M.J. Velgakis. Hyperscaling in the three–dimensional Ising model. Phys. Rev. \underline{B27}, 2839–2854 (1983).

[44] J.L. Gammel, J. Nuttall, D.C. Power. To appear.

[45] M.J. George, J.J. Rehr. Two–series approach to partial differential approximants: Three–dimensional Ising models. Phys. Rev. Lett. \underline{53}, 2063–2066 (1984).

[46] J.–H. Chen, M.E. Fisher. To appear in Journal de Physique.

[47] T. Guttmann. In preparation.